I am trying to write a general conic form equation of a circle usng this information given in the problem:
Ends of diameter: (2,0) and (-18,8)

Answer by lwsshak3(11628)

I am trying to write a general conic form equation of a circle usng this information given in the problem:
Ends of diameter: (2,0) and (-18,8)
**
The procedure used to solve this problem: Find coordinates of midpoint of diameter which are the coordinates of the center. Then use coordinates of one of the given points to find the radius.
..
Standard form of equation for a circle: (x-h)+(y-k)^2=r^2, (h,k)=(x,y) coordinates of center, r=radius
midpoint of diameter:
(-18+2)/2, (8+2)/2=-16/2,8/2=(-8,4)
center: (-8,4)
plug in coordinates of center (-8,4) and given point (2,0), and solve for r.
(2+8)^2+(0-4)^2=r^2
(10)^2+(-4)^2=r^2
100+16=r^2
r^2=116
r≈√116≈10.77
Equation of given circle:
(x+8)^2+(y-4)^2=116